OpenStudy (anonymous):

The length of a pole is to be 6.5 ft.The actual length of the pole may vary by as much as 0.02ft. Write an absolute value inequality for the possible lengths of the pole.

OpenStudy (anonymous):

\[x=6.5\pm0.02\]

OpenStudy (anonymous):

\[\huge |x-6.5|\leq 0.02 \] this is the same thing as saying: \[\huge 6.48\le x\le 6.52 \]